Leonard decides to create 5.0 g of the artificial element neptunium 236. The element is radioactive with a half life of 1.5 x 105 years.


a.     What is meant by an “artificial element”?  (1)



b.     Calculate the initial activity of the neptunium 236 sample. The number of nuclei in 236 g of neptunium are 6.02 x 1023. (3)



c.     Neptunium 236 has three possible decay modes which are shown in the table below. The products of each type of decay are also shown.








1
Expert's answer
2021-02-25T11:24:44-0500

a) A synthetic element is one of 24 chemical elements that do not occur naturally on Earth: they have been created by human manipulation of fundamental particles in a nuclear reactor, a particle accelerator, or the explosion of an atomic bomb; thus, they are called "synthetic", "artificial", or "man-made".


b)


"A=\frac{\ln{2}\cdot 5\cdot 6.02\cdot10^{23}}{236\cdot 1.5\cdot10^5}=5.9\cdot10^{16}\ Bq"